OpenStudy (anonymous):

Please Help! Show me how its done please! Find the sum of the first 8 terms of the sequence. 1, -3, -7, -11, ...

OpenStudy (amistre64):

sooo, you need help adding? or is it that you really need help in determining all the 8 terms?

OpenStudy (amistre64):

since the numbers are getting smaller, or going negative, ... and they seem fairly close together. We can try to see if they are subtracting a like amount each time.

OpenStudy (notamathgenius):

Fine out what the term is going up by. I see -4

OpenStudy (anonymous):

I need help determining all the 8 terms, this is pre cal and it really confuses me. I'm not sure what formula to use. The Arithmetic Sequence formula?

OpenStudy (amistre64):

formulas are only half the battle. seeing a pattern is the key once you see that pattern, the formula is known by default

OpenStudy (anonymous):

So that is it? Its isn't any of the formulas they gave me?

OpenStudy (amistre64):

if you subtract 4 from the last term given, what do you get?

OpenStudy (anonymous):

-15, -19, -23

OpenStudy (amistre64):

.... or if you want to create an expression to use in summation thats fine too \[\sum_{n=a}^b~a_1+d(n-1)\] \[\sum_{n=a}^b~a_1+\sum_{n=a}^bd(n-1)\] \[a_1(b-a)+d\sum_{n=a}^bn-1\] \[a_1(b-a)+d\sum_{n=a}^bn-d\sum_{n=a}^b1\] \[a_1(b-a)+d\sum_{n=a}^bn-d(b-a)\] stuff like that

OpenStudy (amistre64):

but yes ..., -15, -19, -23, -27 fill in all 8 terms

OpenStudy (amistre64):

\[\frac{n}{2}(a_1+a_8)=sum\]

OpenStudy (anonymous):

ok got it! thanks!

OpenStudy (amistre64):

youre welcome, good luck :)